### v1

# Solving Stochastic Orienteering Problems with Chance Constraints Using a GNN Powered Monte Carlo Tree Search

Abstract
Leveraging the power of a graph neural network (GNN) with message passing, we present a Monte Carlo Tree Search (MCTS) method to solve stochastic orienteering problems with chance constraints. While adhering to an assigned travel budget the algorithm seeks to maximize collected reward while incurring stochastic travel costs. In this context, the acceptable probability of exceeding the assigned budget is expressed as a chance constraint. Our MCTS solution is an online and anytime algorithm alternating planning and execution that determines the next vertex to visit by continuously monitoring the remaining travel budget. The novelty of our work is that the rollout phase in the MCTS framework is implemented using a message passing GNN, predicting both the utility and failure probability of each available action. This allows to enormously expedite the search process. Our experimental evaluation shows that with the proposed method and architecture we manage to efficiently solve complex problem instances while incurring in moderate losses in terms of collected reward. Moreover, we demonstrate how the approach is capable of generalizing beyond the characteristics of the training dataset. The paper’s website, open-source code, and supplementary documentation can be found at ucmercedrobotics.github.io/gnn-sop.
I Introduction
Orienteering is an APX-hard optimization problem defined on a weighted graph, $G$, where each vertex has a reward and each edge has a non-negative cost [7]. The goal is to plan a path between designated start and end vertices to maximize the total reward collected from visited vertices while staying within a budget, $B$, that limits the total path length. This budget is often considered in terms of time, power, or distance that can be traveled (see Figure 1). Unlike the traveling salesman problem (TSP), a typical solution to an orienteering problem instance does not visit all nodes. Instances of the orienteering problem can model various real-world scenarios we encounter daily, such as logistics [16], surveillance [25, 14], ridesharing [12], and precision agriculture [21], among others. Our interest in this problem is driven by its applications in precision agriculture [18, 19, 20, 2], though its range of uses is broad and continually expanding.
Figure 1: The orienteering problem is defined on a graph $G$ where vertices have rewards and edges have non-negative costs. The goal is to find a path from an assigned start vertex (the red node in this example) to a goal vertex (the green node) that maximizes the sum of collected rewards while ensuring the path cost does not exceed a given budget, $B$.
Most research on orienteering has concentrated on the deterministic version, where both vertex rewards and edge costs are known in advance. However, this is not often the case in many practical scenarios. For example, when a robot moves from one location to another, the time or energy required for the movement is often a random variable rather than a deterministic value. This variant is known as the stochastic orienteering problem (SOP). Although SOP has been studied [1], it has received significantly less attention than its deterministic counterpart.
Recently, we proposed a novel solution to the stochastic orienteering problem using Monte Carlo Tree Search (MCTS), aimed at estimating both the collected reward and the probability of exceeding the allocated budget $B$ [2]. Previous work highlighted in [24] and confirmed through our own experiments demonstrates that the rollout phase of MCTS can consume up to 99% of the total planning time. Aiming at accelerating the search with the goal of deploying a planner in applications requiring real time response, in this work we investigate the potential of using a graph neural network (GNN) architecture to train a model capable of predicting rollout outcomes. A key innovation of our approach lies in simultaneously predicting two critical metrics simultaneously needed to solve SOP instances: the accumulated reward and the probability of budget constraint violations. To the best of our knowledge, this dual prediction has not been explored previously.
Accordingly, in this contribution we present our GNN-powered MCTS SOP solver. Our novel contributions as follows:
  • we introduce a single network architecture capable of predicting values for a custom developed rollout procedure that outputs both action quality and failure probabilities;
  • we demonstrate that this approach is competitive with the model that trained it and even optimal solutions, at times;
  • we demonstrate that we can plan and execute complex orienteering solutions at real-time speeds.
The rest of the paper is organized as follows. Selected related work is presented in Section II. SOP is formalized in Section III, and in the same section we briefly define the model that generates our training data. The core of our contribution is in Section IV where we describe the network we developed, as well as the feature attributes we identified to effectively train the networks. Extensive simulations detailing our findings are given in Section V, with conclusions and future work discussed in Section VI.
II Related Literature
Early approaches to solve deterministic and stochastic orienteering problems relied on classical methods, often using heuristics due to the intrinsic computational complexity [8]. Notably, in [22] the authors propose an exact solution based on a mixed integer linear program formulation to solve the stochastic orienteering problem with chance constraints. This approach is however offline, i.e., it computes a solution before execution starts, and does not adapt on the fly. Our recent work [2] instead solves the problem using MCTS, designed to be an online method, whereby the vertices to visit are selected at runtime based on the remaining budget. This method, while competitive, devotes a significant amount of time to the rollout phase in MCTS search.
The practice of embedding policy and or value networks into MCTS has been notably used in [15] and in subsequent papers such as [28]. Even though these papers do not use GNNs, they provide evidence into being able to solve computationally expensive problems by reducing the simulation required to approximate a solution. They leverage the power of MCTS and approximate their rollout phases, among other improvements. Several recent approaches have explored the use of GNN-based solutions to solve NP-hard graph optimization problems. For example, [5, 24] focused on using GNNs to solve TSP and other graph optimization problems. While different from orienteering, solutions to TSP are of interest due to some similarities in the problem formulation. Both of these papers used a message passing framework with varying layer configuration. Importantly, [24] used MCTS to explore a limited horizon rather than the entire state space. In [11] the authors show that a similar implementation can be used to solve orienteering problems. In particular, a GNN-based approach is used in combination with a beam search method as done in [24]. However, none of these approaches address stochastic orienteering. Equally relevant, to the best of our knowledge, no methods combining GNNs and MCTS have been developed to solve the SOP with chance constraints, which is the focus of this contribution.
III Problem Statement and Background
In this section, we formally introduce the stochastic orienteering problem with chance constraints (SOPCC). We then outline our previous solution utilizing MCTS, which motivates the method proposed in this manuscript.
III-A Stochastic Orienteering Problem with Chance Constraints
The deterministic orienteering problem is defined as follows. Let $G=(V,E)$ be a weighted graph with $n$ vertices, where $V$ is the set of vertices and $E$ is the set of edges. Without loss of generality, we assume $G$ is a complete graph. Define $r:V\rightarrow\mathbb{R}$ as the reward function assigning a reward to each vertex, and $c:E\rightarrow\mathbb{R}_{+}$ as the cost function assigning a positive cost to each edge. Let $v_{s},v_{g}\in V$ be the designated start and end vertices, respectively, and let $B>0$ be a fixed budget. Note that we allow $v_{s}=v_{g}$ for cases where the start and end vertices are the same. For a path $P$ in $G$, define $R(P)$ as the sum of the rewards of the vertices along $P$, and $C(P)$ as the sum of the costs of the edges in $P$. The deterministic orienteering problem aims to find
$P^{*}=\arg\max_{P\in\Pi}R(P)\qquad\textrm{s.t.}~{}~{}C(P^{*})\leq B$
where $\Pi$ is the set of simple paths in $G$ starting at $v_{s}$ and ending at $v_{g}$. A simple path is defined as a path that does not revisit any vertex. Given the connectivity assumption of $G$, restricting $\Pi$ to simple paths is not limiting. In the stochastic version of the problem, the cost associated with each edge is not fixed but rather sampled from a continuous random variable with a known probability density function that has strictly positive support. Specifically, for each edge $e\in E$ we assume $c(e)$ is sampled from $d(e)$ where $d(e)$ represents the random variable modeling the cost of traversing edge $e$. In general, different edges are associated with different random variables. In the stochastic case then, for a given path $P$ the path cost $C(P)$ is also a random variable. The constraint on the path cost $C(P)$ must therefore be expressed using a chance constraint, defined formally as follows.
Stochastic Orienteering Problem with Chance Constraints (SOPCC) With the notation introduced above, let $10<P_{f}<$ 1 be an assigned failure bound. The SOPCC seeks to solve the following optimization problem:
$P^{*}$ $=\arg\max_{P\in\Pi}R(P)$
s.t. $\Pr[C(P^{*})>B]\leq P_{f}.$
Remark: since the random variables associated with the edges are assumed to be known, generating random samples for the cost $C(P)$ of a path $P$ is straightforward and does not necessitate computing the probability density function of $C(P)$. This is consistent with the assumptions usually made when using MCTS to solve planning problems [9].
The problem definition models a decision maker aiming at maximizing the reward collected along a path while remaining below the allotted budget. Because the cost of the path $C(P)$ is a random variable, the constraint can be satisfied only in probability, and this leads to the introduction of the chance constraint $\Pr[C(P^{*})>B]\leq P_{f}.$
III-B Monte Carlo Tree Search for SOPCC
In [2] we introduced a new approach based on MCTS to solve the SOPCC. Key to our new solution was the introduction of a novel criterion for tree search dubbed UCTF – Upper Confidence Bound for Tree with Failures. In this subsection we shortly review this solving method. We refer the reader to [2] for a deeper discussion of our method and to [4] for a general introduction to the MCTS methodology. Note that in [2] when deciding which vertex to add to the route, MCTS only considers the $K$-nearest neighbors to the current vertex. In the version implemented in this paper to train the GNN described in the next section, we remove this feature and instead consider all vertices. The reason is that in [2] we reduced the number of possibilities to expedite the online search, while here training is done offline. We can therefore consider a larger search space. As pointed out in [17] (chapter 8) a solution based on MCTS relies on the definition of four steps: selection, expansion, rollout and backup. In the following, we sketch how these can be customized for solving SOPCC.
III-B1 Selection
Selection, also known as tree policy, guides the search from the root to a leaf node based on the value associated with the nodes in the tree. At every level, each child of the current node is evaluated using a metric, and the search then proceeds to the node with the highest value. Universal Confidence bound for Trees (UCT) is commonly used to attribute values to nodes [10] and defines the tree policy. UCT, however, aims at maximizing reward only and is therefore not suitable for use in problems with chance constraints where high reward could lead to violations of the constraint. To overcome this limit, in [2] we introduced UCTF, defined as follows
$\begin{gathered}UCTF(v_{j})=Q[v_{j}](1-F[v_{j}])+z\sqrt{\frac{\text{log}(N[v_{p}])}{N[v_{j}]}}\end{gathered}$ (1)
where the term $(1-F[v_{j}])$ is new. In Eq. (1), $Q[v_{i}]$ is the estimate of how much reward will be collected by adding $v_{j}$ to the route, $F[v_{j}]$ is a failure probability estimate for a path going through $v_{j}$, $N[v_{j}]$ is the number of times $v_{j}$ has been explored and $N[v_{p}]$ is the number of times its parent has been explored. Critically, in our original approach both $Q$ and $F$ are estimated via rollout, while in this paper we will train a GNN to quickly predict these values. The term $(1-F[v_{j}])$ penalizes nodes with high estimates for the failure value. As in the classic MCTS method, the term $z$ balance exploration with exploitation.
III-B2 Expansion
Expansion is the process of adding a node to the tree. In our implementation, this is implicitly obtained by assigning a UCTF value of infinity to vertices that are still unexplored (i.e., for which $N[v_{j}]$ is 0). This way, we ensure all children nodes of parent node are explored at least once before one of the siblings is selected again.
III-B3 Rollout
Once a node is added, rollout (also called simulation) is used to estimate the amount of reward that will be collected by going through that node as well as its failure probability. This is done by running a baseline, handcrafted, heuristic policy multiple times. The $F$ value is estimated as the ratio of failed runs over total runs of the heuristic policy, while the $Q$ value is estimated as the average of the returned value limited to the successful runs. A critical aspect of this step that is the necessity of running the heuristic multiple times to numerically estimate $F$. In [2], the heuristic is a mix of random exploration and greedy search.
III-B4 Backup
The backup step in [2] is unique because it propagates back not only the $Q$ value, but also the $F$ value. This is because for each action we must estimate not only the anticipated reward, $Q$, but also the probability, $F$, that the action will eventually result in a budget violation. To this end, two cases are considered by comparing the $F$ value of the parent with the $F$ value returned by the rollout. If the $F$ value for the parent node $v_{p}$ is $\leq P_{f}$, we check if the $F$ value of the newly expanded node is also below $P_{f}$. If it is, we then compare the reward estimate, $Q$. If the child node has a larger estimate with adding the reward of the parent node, we update the parent estimate. If not, we move on. In the case where the parent node has an $F>P_{f}$, we check if the child node $F$ estimate is less than the parent’s. If it is, we update the parent’s $F$ and $Q$ with $F$ being the child’s $F$ value and $Q$ being the child’s $Q$ plus the true, known reward of the parent node (see [2] for additional details).
IV Methodology
IV-A MCTS Value and Failure Network
The main novelty of the method we propose here is a single value/failure network architecture that predicts both the expected value, $Q$, and the failure probability, $F$, of an action, $v_{j}$, to be used in the tree policy UCTF formula from Eq. (1). Therefore, after appropriate training, the MCTS rollout phase with multiple simulations to estimate both $Q$ and $F$ is replaced by a single forward pass on each network, thus dramatically cutting the computation time. As a result, as explained later, when a node is considered for expansion all descendants are generated and evaluated at once. This is different from the classic MCTS approach where node descendants are added and evaluated one by one. This feature greatly extends the set of actions explored while not requiring more computational time.
IV-B Message Passing Neural Network
Given that orienteering problem instances are defined over graphs, a message passing neural network (MPNN) [27, 26] is the natural choice to implement the value and failure networks. MPNNs are graph neural networks in which, at every iteration of learning, nodes of the graph share information with neighboring nodes to generate an $D$-dimensional embedding for each node or for the entire graph. In a typical message passing framework, the message update function encompasses node attributes as well as edge attributes, both of which are critical in the SOPCC problem space. The following equations show the general relationship between an embedding and a message within the framework [6], formulated in terms of the aggregation operation given by Eq. (2) (where the set $\mathcal{N}(v)$ is the set of nodes neighbors of $v$) and the update operation given by Eq. (3).
$m_{v}^{t+1}$ $=\sum_{w\in N(v)}M_{t}(h_{v}^{t},h^{t}_{w},e_{vw})$ (2)
$h_{v}^{t+1}$ $=U_{t}(h_{v}^{t},m_{v}^{t+1})$ (3)
Aggregation defines that a message $m_{v}^{t+1}$ for node $v$ at time $t+1$ is obtained by combining the embedding $h_{v}^{t}$ with all of its neighbors’ embeddings, $h_{w}^{t}$, at time $t$, as well as the attributes $e_{vw}$ of the edges connecting them. The update operation then creates the embedding $h_{v}^{t+1}$ for vertex $v$ at time $t+1$ by combining its previous embedding at time $t$ with the message at time $t+1$ obtained through aggregation. In both operations it is assumed that the data is transformed by a function, i.e., $M_{t}$ for aggregation and $U_{t}$ for update. In our implementation $M_{t}$ is defined as a matrix multiplication between edge attribute linear transformations and node embeddings. This was defined in [6] as
$M_{t}(h_{v}^{t},h^{t}_{w},e_{vw})=A_{e_{vw}}h_{w}^{t}$
The update function $U_{t}$ is defined as a Gated Recurrent Unit (GRU) layer introduced in [3]. Note that [6] states that the GRU will take as input $h_{v}^{t}$ and $h_{v}^{0}$ at every iteration, but we opt for the network from [13] that instead takes in $h_{v}^{t}$ and $h_{v}^{t-1}$. This design choice was made after experimentally observing that the latter gives better results than the former. Finally, we pass the output from the message passing layers into a multi-layer perceptron (MLP) to obtain the $Q$ and $F$ values. In our implementation, we use three message passing layers, i.e., the maximum value for $t$ is three. Critical to the approach are the node attributes to be used for $h$ at time 00, as well as the edge attributes $e_{vw}$ prior to linearization. These attributes will be defined in the remainder of this section. In our use case, we create two different networks using this architecture, one for $Q$ and one for $F$ [^1] The complete GNN architecture as well as relevant training documentation can be found at ucmercedrobotics.github.io/gnn-sop.. We have two separate networks since the MCTS training data is best captured by two different activation functions for $Q$ and $F$, We select a linear activation function for the $Q$ value and a sigmoid for $F$. The reasoning is that $Q$ predicts a positive, but potentially unbounded scalar value while $F$ predicts a probability between 0 and 1. We also hypothesized that the networks would train differently on both types of labels, so we opted to separate them. Indeed, preliminary experiments showed that our loss was reduced when separating the models completely instead of having one model with two output activation functions or two output MLPs. We selected mean squared error (MSE) as the loss function for both $Q$ and $F$ outputs, as the task at hand is regression.
IV-C Attributes
For each node in the graph, we introduce an eight dimensional vector of attributes that will be used to define $h_{v}$ Eq. (2) for $t=0$. Each node has the following attributes:
  • a binary value indicating if the vertex has been visited;
  • its $x$ and $y$ coordinates in the plane;
  • its reward;
  • the remaining budget;
  • a binary value indicating if the node is the start vertex;
  • a binary value indicating if the node is the end vertex;
  • a binary value indicating if the node is the vertex where the robot is currently positioned at.
Our choice for these attributes starts from [24] where a similar approach was used to solve the traveling salesman problem (TSP). Accordingly, the first three attributes give the model the spatial awareness that proved to work well in TSP. However, since SOPCC has more requirements that TSP, we added additional attributes such as reward and residual budget owing to the fact that the nodes are not all uniformly important and in general the agent will not be able to visit each one. Intuition says that more spatial and temporal information will aid the model in learning the greedy heuristic we described in Section III-B. Indeed, as we will show in Section V-B through an ablation study, these additions were critical in performing comparably to previous methods. As for edge attributes $e_{vw}$, we set the single attribute to the Euclidean distance between nodes $v$ and $w$. Note that the edge attributes do not include any information about the length variability $d(e_{v,w})$. This choice was made after preliminary tests showed no conclusive performance improvement and a reduction in model parameters, allowing for a smaller training dataset and, ultimately, reducing training time.
In selecting our model, we started from the literature discussed in Section II as well as a test case comparison. We trained both a MPNN [6] and a graph attention network (GAT) [23]. Our starting hypothesis was that an attention network might be able to identify more valuable nodes through the attention mechanism. However, it turned out that the performance was a bit worse when training the GAT. To be specific, the average reward computed was roughly 10% worse in our largest model, with only  50ms gains in computational time through the entire MCTS pipeline, inference included. This trade off was, therefore, not substantial enough for selecting the GAT. Consequently, we selected the neural message passing model from [6, 13] for its performance and stability. Notably, the pre-built model from [13] includes a skip level connection from the initial node attributes to post-message passing layers. We removed this skip layer to be consistent with [6] and also because we observed it did not yield better results.
IV-D Training
With the goal of replacing the rollout phase in MCTS with a trained model that can estimate both the utility, $Q$, and failure probabilities, $F$, without simulation, a necessary preliminary step is generating data to train the MPNN. To this end, we started by solving randomly generated instances of the SOPCC problem using the MCTS algorithm described in Section III-B. These training instances are complete graphs with rewards and $(x,y)$ coordinates uniformly and randomly distributed between zero and one (see Figure 2) with edge lengths equal to the Euclidean distance. For training, we generated graphs with 20, 30 and 40 vertices.
Figure 2: Examples of randomly generated training graphs with twenty and thirty vertices.
We trained all of our networks on a budget $B=2$, and a failure probability $P_{f}=0.1$. During training, travel costs along edges were sampled from an exponential distribution with mean equal to the deterministic edge length (same noise model used in our preliminary work [2].) We noticed that any network trained on more than 6,000 problem instances solutions seemingly converged. Importantly, for training we only used data from successful SOPCC solutions, i.e., we excluded cases where the MCTS algorithm produced a solution that violated the budget constraint. We will go over the impact of these decisions in Section V.
V Results
V-A Experiments
We experimentally assessed how the architecture we designed predicts the $Q$ and $F$ values to efficiently solve instances of SOPCC. We compare our GNN-MCTS based method with an exact method based on mixed integer linear program [22] (referred to as MILP in the following), as well as the method we formerly presented in [2], referred to as MCTS-SOPCC. For completeness we also compare the GNN-MCTS against the method described in section III-B that was used for training (referred to as MCTS – recall that while it is based on MCTS-SOPCC it has some slight implementation differences as formerly discussed). Our implementation is written in Python 3.11.5 and we used [13] to implement the network based on pre-existing GNN layers. The test machine has an Intel Core i7-10700F at 2.90GHz, 64GB of RAM, and an Nvidia 1660 SUPER GPU. For the MILP based solution, we used the commercial solver Gurobi. The goal is assessing a value network’s ability to learn heuristics and generalize while solving SOPCC instances of varying complexity. Compared to existing solvers, the strength of the GNN-MCTS solver is expected to be speed. Moreover we expect GNN-MCTS to scale much better as the size of problem instances increase. For both MCTS and MCTS-SOPCC comparison, the number of rollouts $S=100$, i.e., $Q$ and $F$ are estimated averaging 100 the results of 100 independent executions. This value was taken from [2]. Both MCTS and MCTS-SOPCC expand the tree 350 times (a value also taken from [2]. In all cases, we set $P_{f}=0.1$.
Figure 3: Time comparison between the different algorithms. GNN-MCTS took always less than half a second, while MILP times averaged 300 to 400 seconds from graph size 20 to 40 (bars cut to save space).
We start by assessing the time performance of the different algorithms. We display these results in Figure 3 for graphs with different number of vertices, and we display averages over 100 different instances. The most evident result from these experiments is how much faster our algorithm is than those existing. Timing performance is increased upwards of 1000% in most instances.
Timing aside, when solving SOPCC instances one has to consider both the reward collected and the failure rate. Accordingly, we define performance through two variables, average reward obtained per solution, $R$, and average rate of failure, $F$. Obviously, a method that obtains a lot of reward, but largely exceeds the failure rate $P_{f}$ is not practically valuable. Hence, a valuable solution is one that stays within the bounds of the constraints while collecting high cumulative rewards. For this test and the following tests, we reused the same test graphs supplied by [2] for their benchmarking in addition to larger sized graphs. Since we trained on instances with $P_{f}=0.1$, we will compare other methods with the same value. In the experiments, as noted in Table I, we see that GNN-MCTS is at times more conservative and achieves lower failure rates, $F$, while incurring limited reward loss. However, in the case of larger, unseen graph sizes, we note larger failure rates. We will touch more on the implications of network generalization in Section V-C.
Figure 4: A test case of size 20 plotted over our test orchard. The white arrows indicate what our GNN-MCTS solved for this problem as the best path.
In Table II we display the performance ratios between GNN-MCTS, the MCTS model from which our GNN was trained, and the MILP solver. This illustrates how truly fast and effective our model is compared to existing methods. We stayed within 93% or better of our MCTS model, showing that this architecture loses less than 10% performance in these cases. Even more impressive, in some cases we out performed the original model. While performance against the MILP was 60% or better with respect to reward, we can expect improvement with optimal training data, which is much more difficult and time consuming to generate. Figure 5 compares the solutions produced by three of the compared algorithms for an SOPCC instance with 40 vertices.
$P_{f}$ GNN-MCTS MCTS MCTS-SOPCC MILP
Reward Time (s) $F$ Reward Time (s) $F$ Reward Time (s) $F$ Reward Time (s) $F$
[…]
TABLE I: Overall results from benchmarks used in [2] plus even larger graphs averaged over 100 trials. Top rows are for results against previously seen budgets. signify tests run with a model trained against size 40 graphs. & are for results against unseen budgets using same sized model. ! is further explained in V-C.
(a) Graph40 GNN-MCTS solution
(b) Graph40 MCTS solution
(c) Graph40 MILP solution
Figure 5: Proposed solutions for three algorithms graph40 test described in Table I. The green node is start; red is goal. The blue represents nodes of varying reward, darker being higher.
GNN-MCTS / MILP GNN-MCTS / MCTS
$R$ Ratio Time Ratio $R$ Ratio Time Ratio
[…]
TABLE II: Ratios between our solver and baseline solvers for reward and time averaged over 100 trials. See Table III for unseen graph sizes.
It is also interesting to explore how the solver performs as $P_{f}$ varies. While in most cases model performance characterized by $P_{f}$ was as expected, a limitation of our implementation is that the activation function modeling $F$ is a sigmoid. This aspect is further discussed in Section V-C.
Ultimately, we achieved the results of real-time approximation in a stochastic environment. Figure 4 shows the solution produced by GNN-MCTS while solving an instance of SOPCC associated with a sampling problem in precision agriculture. The blue dots represent locations to visit, with the color intensity indicating the reward (darker means higher reward). The robot starts at the green vertex and must end at the red vertex before running out of budget.
V-B Ablation Study
Previous literature showed that including spatial information into the vertices attributes is essential when assessing the ability of GNN based methods to solve problems like TSP [5]. However, since solutions of SOPCC instances are also influenced by rewards and remaining budget as well as start and goal locations as pointed out in Section IV-C in our implementation, we correspondingly extended the set of attributes. Therefore, this subsection, through an ablation study, we assess the impact of the temporal and spatial attribute additions, novel to this paper, of the one-hot encoding of the currently visited vertex and the start/goal vertices.
Figure 6: Ablation study substantiating that both temporal and spatial node attribute additions are needed for the value network to perform effectively as removing either one implies decrease in reward collected (left) or increase in failure rate (right). Data averaged over 100 trials with $P_{f}=0.1$.
In Figure 6, we see that when we remove any of these additional attributes, we incur in a decrease in average reward. In the case of start/end vertex removal, we saw nearly a 50% reward loss. However, the most significant difference is failure rate, which significantly violates the $P_{f}$ constraint making this an infeasible solution. The same figure also shows that without the current vertex attribute we do not violate $P_{f}$, but the average total reward is worse than the original model. While this solution is more conservative, having a lower $F$ rate, our original model doesn’t violate $P_{f}$ while collecting – on average – more reward. For node level classification using GNNs, these findings are relatively important in that GNNs possess the capability to understand time and space with respect to a given problem. Instead of embedding that information elsewhere in the algorithm, we can instead task the GNN with understanding explicitly defined spatiotemporal data.
V-C Generalization
As with every learning based approach, the quality of the solution is influenced by the nature of the dataset used for training. In our work, with the explicit goal of assessing the ability to learn and generalize, during training we kept $P_{f}=0.1$ and $B=2$, and used the same underlying graph topology, i.e., complete graphs with vertices uniformly scattered in the unit square. With these premises, it is essential to assess how the proposed architecture can generalize to instances characterized by new parameters. We experimented generalization from two different angles: generalization of graph sizes and unseen budgets. The ability to generalize with respect to these two parameters is not only important for robustness, but also to expedite training, especially when it comes to graph size, as larger instances require more time to generate solutions with the MCTS algorithm used for training. As pointed out in Section IV-D, during training we only considered graphs with 20, 30, and 40 vertices.
For generalizing with respect to size, we saw in Table I, within a certain size range, equally competitive performance. For example, we pushed our model to the high end of what we have seen in SOPCC literature (e.g., up to 70 vertices) and found that the generalization is quite effective in terms of reward performance. However, due to the sigmoid problem previously described and well known in literature, for larger problem instances the model requires a tuned $P_{f}$ value. This means that, as displayed in Table I, if one feeds the model with a desired $P_{f}$ value, for unseen problem instances we get solutions with failure probabilities significantly exceeding the limit. This problem can however be remedied by feeding the model with a $P_{f}$ value smaller than the desired one. Note that this hyperparametrization process is necessary also when using the exact method [22] and is not a limitation unique to the method we present. Table III shows how by adjusting the $P_{f}$ value in unseen problem instances GNN-MCTS manages to successfully solve the problem matching the originally desired failure rate.
$P_{f}$ $R$ $F$ MCTS $R$ Ratio
[…]
TABLE III: When hyperparameterizing $P_{f}$ on unseen graph sizes, we see the model performance improve to near ground truth, presumably due to the sigmoid activation function.
In the case of budget generalization, we can see in Table I that our method was able to maintain similar time and performance ratios, shown in Table II, for unseen budgets. Given that solving a SOP with a lower budget is more difficult, it makes sense that the more budget that is given, the simpler the task becomes for the model. However, we see our algorithm with noticeably lower $F$ without $R$ sacrifice implying – again – that our method is more conservative without much performance impact.
VI Conclusions and Future Work
In this paper, we have presented a new solution to the SOPCC combining MCTS with a message passing graph neural network. This is accomplished by training a neural network that can replace the time consuming rollout phase in MCTS. Key to the idea of solving SOPCC using MCTS introduced in [2], the trained model simultaneously predicts both the anticipated collected reward, $Q$, as well as the predicted failure probability, $F$ – a feature not found in previous works using value networks with reinforcement learning. Another advantage of the methodology we presented is that during the expansion phase, a single forward pass of the network makes predictions for all children at once, instead of considering one child at a time. This allows the algorithm to efficiently explore a larger fraction of the search space.
Our method has been trained using SOPCC solutions computed by a slightly modified version of the algorithm presented in [2]. Extensive simulations compared this new solution with the training method, our former MCTS algorithm presented in [2], and an exact solution based on MILP. Our findings show that the proposed method is orders of magnitude faster while meeting the failure constraint and retaining a significant fraction of the collected reward, though in some cases this requires tuning some hyperparameters. Our ablation study also corroborates that the vertices attributes we selected for the embedding are critical for the success of the method.
There are numerous venues for current and future work. First, it would be interesting to train the model with solutions to SOPCC instances produced by different heuristics or, possibly, an ensemble of heuristics and assess if this improves the overall performance, especially when it comes to generalize to large problem instances. Next, it will also be interesting to explore whether different vertex or edge attributes would improve the predictive capability of the currently used network. Along the same lines, we will also explore whether altering the structure of the final layers of the network (e.g., substituting the sigmoids) would result in better results. Finally, we are currently implementing the planner in ROS 2 and integrating it into our current autonomous navigation stack for precision agriculture and assessing how it can be used to solve in real time complex data collection problems in the field.